OpenStudy (natasha18):

Given the functions, f(x) = x - 8 and g(x) = x2 + x - 1, perform the indicated operation. When applicable, state the domain restriction. (fg)(x) x3 - 7x2 - 9x + 8 x3 - 9x2 - 9x + 8 x3 - 9x2 - 9x - 8 x3 - 7x2 - 9x - 8

OpenStudy (natasha18):

I have no idea how to figure this thing out...

OpenStudy (kg1975):

Try Google! XD

OpenStudy (natasha18):

I tried but nothing came up

OpenStudy (kingnoob):

It's asking you to multiply the functions - \(\large (x-8)*(x^2+x-1) \)

OpenStudy (skullpatrol):

$$\huge (fg)(x)=f(x)*g(x)$$

OpenStudy (natasha18):

Whenever I multiply I don't get any of my answer options

OpenStudy (sshayer):

\[(x-8) \times (x^2+x-1)=x^3+x^2-x-8x^2-8x+8=?\]

OpenStudy (natasha18):

im not sure where you're getting all the extra x's from

OpenStudy (natasha18):

I can see the x ubed and the x squared

OpenStudy (skullpatrol):

$$a(b+c+d)=ab+ac+ad$$

OpenStudy (skullpatrol):

Let a = x-8

OpenStudy (sshayer):



OpenStudy (skullpatrol):

Then use$$ (b+c)a=ba+ca$$